OpenStudy (tabby):

(3-the square root of 7)(3+square root of 7)

OpenStudy (anonymous):

this is in the form (x-y)(x+y)=x^2-y^2. (3-sqrt7)(3+sqrt7)=9-7=2

OpenStudy (tabby):

that is not one of my choices, but thanks.